### v1

# A Scalable Unsupervised Framework for multi-aspect labeling of Multilingual and Multi-Domain Review Data

Abstract
Effectively analyzing online review data is essential across industries. However, many existing studies are limited to specific domains and languages or depend on supervised learning approaches that require large-scale labeled datasets. To address these limitations, we propose a multilingual, scalable, and unsupervised framework for cross-domain aspect detection. This framework is designed for multi-aspect labeling of multilingual and multi-domain review data. In this study, we apply automatic labeling to Korean and English review datasets spanning various domains and assess the quality of the generated labels through extensive experiments. Aspect category candidates are first extracted through clustering, and each review is then represented as an aspect-aware embedding vector using negative sampling. To evaluate the framework, we conduct multi-aspect labeling and fine-tune several pretrained language models to measure the effectiveness of the automatically generated labels. Results show that these models achieve high performance, demonstrating that the labels are suitable for training. Furthermore, comparisons with publicly available large language models highlight the framework’s superior consistency and scalability when processing large-scale data. A human evaluation also confirms that the quality of the automatic labels is comparable to those created manually. This study demonstrates the potential of a robust multi-aspect labeling approach that overcomes limitations of supervised methods and is adaptable to multilingual, multi-domain environments. Future research will explore automatic review summarization and the integration of artificial intelligence agents to further improve the efficiency and depth of review analysis.
keywords:
Multi-Aspect Labeling , Unsupervised learning , Domain-Agnostic Framework , Multilingual Review Analysis , Automatic Labeling
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
Proposes a scalable framework for multi-aspect category labeling without supervision.
Discovers semantically similar aspect terms and clusters them without predefined labels.
Learns aspect semantics using multi-head attention, negative sampling, and clustering.
Assigns aspect categories using prompts in zero- and few-shot settings with language models.
Achieves strong performance across domains and languages, including English and Korean.
1 Introduction
User review data contains direct feedback from consumers about products and services, and can be used as an important source of information for business decision-making, marketing strategy development, and brand reputation management. However, most review data exists in unstructured text format, and typical sentiment polarity analysis or star rating–based analysis has the limitation of not sufficiently capturing the detailed attributes embedded in the reviews. As an alternative to overcome these limitations, Aspect-based Analysis has been gaining attention. Aspect-based Analysis can be widely used in various fields such as consumer feedback analysis, public policy monitoring, and social media analysis (Khabiri et al., 2021; Qiu et al., 2011), and has the advantage of enabling detailed understanding of how each aspect is evaluated in text.
However, even Aspect-based Analysis has several limitations when applied to user review data, which can be broadly categorized into four aspects. First, most studies on Aspect-based Analysis focus on specific domains such as electronics or restaurant reviews (Zhao et al., 2020), making it difficult to ensure that the proposed methods show the same performance and validity in real service environments where a wider range of domains must be addressed. In other words, studies conducted within limited domains do not sufficiently demonstrate multi-domain applicability, which restricts their practical use across industries. Second, Aspect-based Analysis research tends to focus on quantitative metrics such as accuracy and F1-score (Zhang et al., 2024), and lacks qualitative research verifying whether the aspects extracted by the model are semantically valid and practically effective in the real world. As a result, it becomes difficult to determine whether the extracted aspects are trustworthy enough to be used for problem diagnosis or product improvement in practice, which again acts as a limitation in applying the methods in industry. Third, most Aspect-based Analysis research adopts approaches that fine-tune labeled data or assume specific data formats, which leads to high labeling costs and necessitates format conversion (Mukherjee et al., 2021). When relying solely on labeled data, it becomes difficult to accurately capture aspects from newly appearing review texts that are not labeled, leading to limitations in responding in real time in actual service environments. Lastly, another issue is that most Aspect-based Analysis studies are conducted mainly on English data. Since large-scale review data is accumulated in various languages in global service environments, models trained only in a specific language face limitations in expanding to multilingual data. As a result, reviews written in newly emerging or underrepresented languages may not be sufficiently analyzed, leading to missed opportunities for satisfaction analysis and service improvement in various markets.
Therefore, this study aims to overcome the limitations of existing Aspect-based Analysis and proposes MUSCAD (Multilingual and Scalable framework for Cross-domain Aspect Detection), an unsupervised multi-aspect labeling framework for multilingual and multi-domain review data. The core ideas of MUSCAD are as follows:
  1. (a)
    Without large-scale labeling costs or domain-specific formatting, K-means clustering is applied directly to unlabeled review texts to automatically group semantically similar words and generate initial aspect category candidates.
  2. (b)
    To extract aspect-aware embedding vectors that reflect contextual dependencies even across diverse sentence structures, Multi-Head Attention is applied to finely learn key information within sentences, and Max-Margin Loss is used to sharpen the boundaries between aspects.
  3. (c)
    Finally, to ensure scalability across multiple domains and languages, the aspect category candidates extracted through K-means are refined by leveraging both domain expert knowledge and GPT-based models to perform-domain and language-specific naming with improved readability.
To evaluate the effectiveness of MUSCAD, we first perform large-scale automatic labeling on hotel, food, and beauty review datasets written in Korean and English. Next, to verify whether the labels generated by the unsupervised method are practically usable in downstream tasks or supervised learning, we fine-tune classification models and compare performance in terms of accuracy (F1-score). Additionally, we prompt large language models to perform multi-aspect category labeling on the same dataset, and compare the results with the automatically generated labels to verify how consistent and accurate they are compared to LLMs. Lastly, to qualitatively confirm the semantic validity of the extracted labels, we conduct Human Evaluation and compare the results with expert-annotated labels. Through this series of experiments, we aim to demonstrate that the proposed framework can effectively operate in multilingual and multi-domain environments and generate automatically labeled data of sufficient quality for real-world use.
The main contributions of this paper are summarized as follows:
  • Proposal of an unsupervised multi-aspect labeling framework: We propose a new framework that automatically extracts and groups multi-aspect category without manual labeling.
  • Verification of multilingual and multi-domain scalability: We conduct experiments on Korean and English review datasets and empirically validate the framework’s applicability to diverse domains.
  • Quantitative and qualitative performance evaluation: We assess the reliability of the automatically labeled data through comparison with pre-trained classifiers and conduct qualitative evaluation through human annotation.
  • Provision of multi-aspect labeled datasets: We construct high-quality automatically labeled datasets and provide them for use in various downstream tasks through fine-tuning.
2 Related Work
This section reviews previous research on Aspect-based Analysis and examines studies related to its applicability across multiple languages and domains. It also covers recent trends in attention mechanisms and their applications.
2.1 Aspect-based Analysis
Research focusing on the analysis of specific aspects in various types of text data has been actively conducted. This technique is utilized in diverse fields such as consumer review analysis, public opinion monitoring, stock market trend prediction (Lee et al., 2018b, a; Park and Kim, 2019), and online marketing strategy development (Mukherjee et al., 2020; Angelidis et al., 2021).
Early studies mainly relied on manually defined rule-based approaches, applying topic modeling methods such as Latent Dirichlet Allocation (LDA) to extract aspect terms from text (Khabiri et al., 2021). Qiu et al. (2011) proposed extracting aspects using part-of-speech tagging and word frequency data. Another method involved setting preprocessing rules, performing frequency analysis on text data, and manually assigning keywords to predefined categories based on domain knowledge to evaluate keyword importance per attribute (Shofwatalloh et al., 2024). Zhao et al. (2020) proposed a method to extract aspect terms and opinion terms simultaneously, but since it mainly focused on identifying individual elements, it had limitations in effectively extracting aspect-opinion pairs. These approaches depend on predefined rules, making them difficult to apply to various domains, and because aspect categories are selected based on domain knowledge, they have limitations in reflecting domain-specific terms. Moreover, since most analyses are limited to nouns, it is difficult to conduct sentiment analysis that reflects context.
To overcome these limitations, supervised learning–based approaches have been introduced. Early studies explored the use of Conditional Random Fields (CRFs) to simultaneously extract structural features of sentences and polarity opinions (Li et al., 2010). More recently, deep neural network–based approaches have shown promising performance. Wang et al. (2016) applied an attention-based Long Short-Term Memory (LSTM) model for aspect-level sentiment classification, allowing the model to focus on important words in the sentence based on the input aspect. Fan et al. (2019) defined a method called Target-oriented Opinion Words Extraction (TOWE) using Inward-Outward LSTM to extract opinion words related to specific targets. Additionally, the RNCRF model, which combines Recurrent Neural Networks (RNNs) and CRFs, was introduced to learn the relationship between aspects and opinions, and a joint learning method was proposed by optimizing CRF parameters (Wang et al., 2016). Topic-level sentiment analysis research has also been conducted. Approaches using Online Latent Semantic Indexing and LSTM-based models have been proposed to classify sentiment for specific topics in sentences (Pathak et al., 2021). Mikolov et al. (2013) conducted research on learning word representations using neural network–based language models. However, supervised deep learning models require a large amount of labeled data, which involves high cost and time for data construction. Moreover, models trained in a specific domain are difficult to apply directly to other applications.
To address these limitations, recent studies have drawn attention to unsupervised learning–based approaches, which aim to perform effective aspect extraction and sentiment analysis without labeled data. Local topic models at the sentence level have been proposed to automatically infer aspects and generate seed sets of sentiment words without manual annotation (Brody and Elhadad, 2010). Mukherjee and Liu (2012) proposed a semi-supervised topic model that jointly models aspects and sentiment. Zhao et al. (2010) combined topic modeling and maximum entropy (MaxEnt) models to separate aspect and opinion words using a semi-supervised approach. More recently, unsupervised aspect term extraction using attention-based Deep Neural Networks (DNNs) has been proposed (He et al., 2017). These approaches have the advantage of automatically clustering semantically similar aspects without labeled data. However, existing unsupervised methods lack sufficient verification of cross-domain generalization performance, and due to instability during training, they often suffer from low reproducibility and inconsistent performance.
Most previous studies have been conducted in single-language settings. For example, in Aspect-Based Sentiment Analysis using Korean data, Park and Shin (2020) explored Attribute Category Sentiment Classification (ACSC) using Bidirectional Encoder Representations from Transformers (BERT). They replaced the [CLS] token with output vectors from tokens related to attribute categories and analyzed performance differences according to input configurations (QA/NLI). In a study using beauty commerce reviews, Importance-Performance Analysis (IPA) and DistilBERT-based sentiment analysis were used to identify areas for product improvement (Shofwatalloh et al., 2024). Another study analyzed sentiment for specific evaluation categories such as food, price, service, and atmosphere using Korean restaurant reviews and proposed a model for predicting detailed ratings (So and Shin, 2020). Lee and Ahn (2022) applied attribute-based sentiment analysis using BERT in a movie recommendation system to classify reviews by attributes (e.g., direction, actors, story) and develop a Movie Multi-Criteria recommendation model reflecting user preferences. These studies demonstrate that Aspect-Based Sentiment Analysis enables more fine-grained evaluation of products and services and can enhance the performance of recommendation systems. However, they still lack sufficient verification of generalization across domains and studies considering multilingual applicability remain scarce.
2.2 Attention Mechanism in Aspect-based Analysis
Recently, attention mechanisms have been widely used in Natural Language Processing (NLP), significantly enhancing the performance of deep neural network models. In particular, pre-trained models such as BERT, Robustly Optimized BERT Pretraining Approach (RoBERTa) effectively capture contextual information through attention, achieving state-of-the-art performance in various NLP tasks. Attention mechanisms learn by focusing on important information within the text. For example, in sequence tagging tasks, they highlight specific words within context (Li et al., 2021), and in sequence-to-sequence learning, they help maintain core information during sentence transformation (Liu et al., 2019). In sentiment analysis, attention helps models perform more precise predictions by focusing on words related to specific aspects (Wang et al., 2021), and in machine translation, it contributes to improving semantic alignment between source and target languages (Zhang et al., 2023). Many researchers have explored the use of attention mechanisms in supervised models for effective Aspect-Based Sentiment Analysis (ABSA) (Xu et al., 2019). Additionally, unsupervised attention-based models have been proposed to extract aspects from text without large-scale labeled data (He et al., 2017). However, most of these existing studies focus on single languages and specific domains, and there is still a lack of research verifying generalization performance in multilingual and multi-domain environments.
3 Dataset
In this study, we collected review data written in Korean and English to evaluate the performance of our framework in multilingual and multi-domain settings. Specifically, we gathered hotel and food reviews in Korean and beauty reviews in English, applying different preprocessing pipelines according to the linguistic characteristics of each language.
For Korean review data, we used the Korean String Processing Suite (KSS) [^1] https://github.com/hyunwoongko/kss to segment the text into individual sentences for sentence-level analysis. As a result, we obtained 313,333 sentence-level entries for hotel reviews and 288,874 for food reviews. After removing unnecessary special characters and extra spaces, we refined the text using the Open Korean Text (OKT) morphological analyzer from Konlpy (Park and Cho, 2014). Since Korean features various grammatical transformations including verb and adjective inflections, we lemmatized these parts of speech and removed case particles (e.g., Korean case particles such as i/ga (subject), eul/reul (object), e (to/at), eseo (from/in)) to ensure analytical consistency. For example, the Korean expressions “맛있었다” (delicious + past tense) and “맛있을 것이다” (delicious + future tense) were both normalized to “맛있다” (delicious). This preprocessing step was critical for reducing variability caused by Korean-specific tense and ending changes that could interfere with model training. We also customized the OKT user dictionary by adding neologisms, slang, and domain-specific proper nouns. This helped prevent incorrect token splits of words not included in the default dictionary and improved the performance of the analyzer. Both preprocessed and original sentences were retained to enable flexible follow-up analyses without data loss.
For English review data, we used the sent_tokenize function from the NLTK library [^2] https://github.com/nltk/nltk_data to segment the text into sentences. All text was lowercased, and special characters were removed. Then, stopwords were filtered out to eliminate words irrelevant to the analysis. Since English features relatively less morphological variation than Korean, we only applied stopword removal and word normalization without additional lemmatization steps. As a result of these preprocessing steps, we constructed a dataset covering three domains: hotel, food, and beauty, as shown in Table 1.
Domain Language Number of Reviews Number of Sentences
Beauty English 50,000 255,110
Hotel Korean 225,412 313,333
Food Korean 192,484 288,874
Table 1: Overview of the Hotel, Food, and Beauty dataset with review and sentence counts for each domain in English and Korean.
4 Framework Structure
MUSCAD is designed to automatically extract aspect terms and aspect categories from sentences and to learn their interrelationships. Its architecture incorporates K-means clustering, aspect representation learning through self-attention and multi-head attention, and aspect classification optimization using max-margin loss. The overall structure of the proposed framework is shown in Fig. 1.
Figure 1: MUSCAD  Framework
Additionally, the extracted aspect terms and categories can be used to automatically assign aspect category labels to new review data, enabling more precise aspect-aware labeling. Section 4 describes each sub-module in detail.
4.1 Aspect Category Candidate Embedding
In the input processing stage, we perform preprocessing to effectively extract aspect terms from natural language sentences. In this study, we apply a method that learns word embeddings without predefined vocabularies and automatically derives initial aspect category candidates through K-means clustering. This enables scalable extraction of aspect terms and aspect categories that are not domain-dependent.
4.1.1 Embedding Vector Extraction
We train word embeddings reflecting semantic similarity using the Word2Vec model (Mikolov et al., 2013). Specifically, we apply the Continuous Bag of Words (CBOW) approach, which predicts a center word based on its surrounding context words. For example, in the sentence “The hotel has great service and staff”, if “service” is the center word and the window size is 1, the context words are “great” and “and”.
Each context word is represented as a one-hot encoded vector $x_{i}$, which is multiplied by a weight matrix to obtain an embedding vector. The average of these context embeddings becomes the context vector, which is passed through the output weights and a Softmax function to predict the probability of the center word. By minimizing the loss function during training, we obtain the embedding matrix $E_{i}$, which contains embedding vectors for all words.
4.1.2 Unsupervised Generation of Aspect Category Candidate
In the unsupervised aspect extraction stage, since specific aspect terms or categories are not predefined, we must effectively group semantically similar words. We propose applying K-means clustering to the word embeddings generated in Section 4.1.1 to automatically generate initial aspect category candidates, which are later refined into final aspect categories.
K-means clustering is a representative unsupervised learning technique for grouping data points, and it can automatically cluster word embedding vectors based on their semantic similarity. The result of clustering the data using the K-means algorithm can be represented as shown in Eq. 1.
$C_{k}=\{e_{w_{i}}\ |\ \arg\min_{k}\|e_{w_{i}}-c_{k}\|^{2}\},$ (1)
where $e_{i}$ denotes the vector representation of each word, $c_{k}$ is the centroid of the $k^{th}$ cluster, and $C_{k}$ represents the set of words belonging to the $k^{th}$ cluster. Each time a word is reassigned to a new cluster, the cluster centroid $c_{k}$ is updated and then $L_{2}$-normalized to maintain a consistent scale (see Eq. 2).
$\hat{c}_{k}=\frac{c_{k}}{\|c_{k}\|_{2}},~{}~{}where~{}~{}c_{k}=\frac{1}{|C_{k}|}\sum_{e_{w_{i}}\in C_{k}}e_{w_{i}}$ (2)
By repeating this process, we group word vectors into semantically coherent clusters and derive optimal initial aspect category candidates based on the centroid vectors. This allows the model to learn relationships between words and aspect categories effectively without predefined vocabularies or labels in multilingual and multi-domain environments.
4.2 Aspect-Aware Representation Learning
To learn the relationship between each word and aspect category and reflect contextual information in the sentence, we train attention mechanisms. Section 4.2 describes how we use self-attention and Multi-Head Attention to extract fine-grained, context-aware embedding vectors.
4.2.1 Multi-Head Attention-Based Sentence Representation
Attention mechanisms learn the relationships between words in a sentence and reflect the importance of each word in context. Unlike earlier studies (He et al., 2017) that relied only on attention, we use Multi-Head Attention to learn sentence representations that capture contextual information.
In this study, the input embedding $E_{i}$ is fed into the Multi-Head Attention module, where trainable weight matrices are applied to compute the Query, Key, and Value representations, followed by the attention operation. The computation process of the Multi-Head Attention is shown in Eq. 3.
$\texttt{MultiHead}(E_{i},E_{i},E_{i})=\texttt{Concat}(\text{head}_{1},..,\text{head}_{h})W^{O},$ (3)
$W^{O}$ is the output weight matrix that projects the concatenated outputs of all attention heads into the final output dimension. Here, $head$ denotes an individual attention head, and each attention head is computed as shown in Eq. 4.
$\text{head}_{i}=\texttt{softmax}\left(\frac{E_{i}W_{i}^{Q}(E_{i}W_{i}^{K})^{T}}{\sqrt{d_{k}}}\right)E_{i}W_{i}^{V},$ (4)
Here, $W^{Q}_{i}$ projects the input embedding $E_{i}$ into the query space in the $i^{th}$ attention head, $W^{K}_{i}$ projects $E_{i}$ into the key space within the same head, and $W^{V}_{i}$ projects $E_{i}$ into the value space. Through this Multi-Head Attention process, the final context vector $z^{o}$ is generated, capturing the relationships between words in the sentence via weighted summation. While previous studies often derived the sentence representation by directly aggregating the outputs of the Multi-Head Attention, our approach refines contextual representation by multiplying the original word embedding $E_{i}$ with the context vector $z^{o}$, resulting in the final sentence representation vector $z_{i}$.
The resulting $z_{i}$ represents the core meaning of the sentence and is utilized in the subsequent learning process to extract aspect-aware embeddings.
4.2.2 Aspect-Aware Embedding Vector
In this study, the aspect-aware embedding vector $r_{i}$ is computed as shown in Eq. 5, by leveraging the aspect embedding matrix $E_{c}$ calculated in Section 4.1 and the probability vector $p_{t}$, which represents the likelihood of the sentence being associated with each aspect category.
$r_{i}=E_{c}^{T}p_{t},~{}where~{}p_{t}=\texttt{softmax}(W_{p}^{T}z_{i}),$ (5)
Here, $W_{p}$ is the weight matrix that maps the sentence representation $z_{i}$ into an aspect distribution. Through this process, the $k$ aspect candidates initially generated by K-means are further refined during training, and the final aspect-aware embedding vector is determined by allowing the sentence representation $z_{i}$ to select the most relevant aspect.
4.2.3 Optimization of Aspect-Aware Representations
In this study, we adopt the Max-Margin Loss to train the aspect-aware embedding vector $r_{i}$ introduced in Section 4.2.2. In neural language models, the negative sampling technique is widely used to encourage the model to increase the similarity between a sentence and its corresponding aspect category, while reducing the similarity with negative samples. During the application of the negative sampling method, the model is guided to learn more generalizable patterns by minimizing the inner product between the sentence vector $z_{i}$, generated in Section 4.2.1, and negative samples, while maximizing the similarity with the most relevant aspect. The negative samples $z_{n}$ are composed of representation vectors of sentences unrelated to the input sentence, which helps prevent overfitting to a specific aspect during training. To achieve this, we define the hinge loss function as shown in Eq. 6.
$\mathcal{L}=\min\sum_{i=1}^{m}\max(0,1-\langle z_{i},r_{i}\rangle+\langle z_{i}^{(n)},r_{i}\rangle),$ (6)
[…]
4.3 Aspect Category Assignment
In this section, we refine the aspect category candidates derived through unsupervised learning into more interpretable aspect categories. Although K-means clustering groups semantically similar words, the resulting clusters are merely collections of related terms and are not directly usable as aspect categories. Therefore, we perform a post-processing step that analyzes the word lists in each cluster and converts them into semantically meaningful aspect categories. For example, a cluster in hotel reviews containing words such as “modern”, “cozy” and “sophisticated” can be assigned to an aspect category representing room ambiance. Similarly, a cluster in food reviews including terms like “salty”, “nutty” and “tender” can be interpreted as referring to the taste of the food. To facilitate this process, we consulted Korean hotel and food review websites as well as international beauty product review sites to construct a set of candidate aspect categories. Then, using GPT-3.5-turbo, we automatically recommended the most appropriate aspect category for each aspect term cluster. Few-shot prompting was applied based on prompts such as the one shown in Fig. 2. Fig. 2 provides an example of an aspect categorization prompt for hotel reviews, while aspect category candidates for the food and beauty domains are shown in the Appendix Fig. A.1.
Aspect Categorization Prompt (Hotel Domain) Refer to the given aspect term list and select the most appropriate aspect category.
Try to match it with the aspect category list below, but if no suitable match is found, you may suggest a new category.
Aspect Category Candidates: Cleanliness, View, Service, Facilities, Room, Pool, Parking, Breakfast, Amenity, Location, Satisfaction
Figure 2: Hotel Aspect Categorization Prompts
5 Experimental Result
This section comprehensively evaluates the performance and practicality of the proposed MUSCAD framework through a series of experiments. First, we empirically determined the optimal number of clusters for each domain during the construction of the automatically labeled dataset using the framework. The number of aspect categories per domain was also determined through expert review. Then, we conducted multi-label classification experiments using the constructed datasets to compare the performance of fine-tuned classification models with that of LLM–based labeling. In addition, a Human Evaluation was conducted to assess whether the automatically generated labels are semantically valid in real-world domain settings. Through this, we verify the reliability and scalability of the proposed framework.
5.1 Experimental Settings
In the automatic labeling process, which is one of the core components of the proposed framework, the optimal number of clusters and word extraction criteria were determined based on the characteristics of each domain. Details of the constructed datasets and the corresponding settings are summarized in Table 2.
Domain Number of Clusters Number of Terms Final Review Count
Beauty 13 150 255,110
Hotel 15 150 263,395
Food 17 150 210,272
Table 2: Domain-specific Clustering and Auto-labeled Dataset Statistics
The number of clusters and the number of extracted terms per cluster were determined through experiments with various hyperparameter settings. Specifically, we varied the number of extracted terms from 50 to 250 in increments of 50 to identify the optimal configuration. Detailed experimental results on the number of terms per aspect cluster are provided in A. In summary, we used 15 clusters for the hotel domain, 17 for the food domain, and 13 for the beauty domain, extracting the top 150 terms from each cluster to form the initial pool of aspect term candidates. The number of heads in the Multi-Head Attention module was set to $h=4$ to effectively capture diverse contextual information within sentences.
To validate the quality of the labeled datasets generated by the framework, we fine-tuned several pre-trained language models and evaluated their classification performance. We randomly undersampled 40,000 samples for each domain. For the Korean datasets (Hotel and Food), we used KoELECTRA (Park, 2020), KR-BERT (Lee et al., 2020), KLUE-BERT (Park et al., 2021), Multilingual BERT (Pires et al., 2019), and XLM-RoBERTa (Goyal et al., 2021). For the English dataset (Beauty), we used Multilingual BERT, RoBERTa (Liu et al., 2020), BERT-base, BERT-large (Devlin et al., 2019), XLM-RoBERTa, and DeBERTa-base (He et al., 2021). All pre-trained language models were fine-tuned using the same hyperparameter settings: a batch size of 16, a maximum token length of 128, and a learning rate of 5e-5. We used the AdamW optimizer and trained each model for 5 epochs. During training, we applied 5-fold cross-validation to ensure stable performance across the dataset and evaluated the models using both Micro-F1 and Macro-F1 scores.
In addition, to compare the performance of fine-tuned models with large language models (LLMs), we conducted multi-aspect labeling using GPT-3.5-turbo and GPT-4o-mini. Few-shot learning was applied in the LLM experiments, and due to API cost constraints, we randomly undersampled 1,000 samples per domain for evaluation. The fine-tuned models for comparison were trained and evaluated on the same datasets using 5-fold cross-validation. For the LLM experiments, batch sizes ranged from 8 to 16, the maximum token length was set to 128, and the learning rate was set between 1e-5 and 3e-5. All experiments were conducted using a NVIDIA GeForce RTX 3090 8GPU environment.
5.2 Finalization of Aspect Categories
The aspect category lists automatically recommended in Section 4 were reviewed and refined by domain experts, resulting in the final categories shown in Table 3.
Domain Count Aspect Category
Beauty 10 Satisfaction, Improvement, Usage method, Scent, Color, Persistence, Purchase, Packaging, Ingredients, Hair
Hotel 7 룸, 서비스, 만족도, 목적, 위치, 뷰, 부대시설 (Translation: Room, Service, Satisfaction, Purpose, Location, View, Facilities)
Food 9 음식, 음식량, 맛, 만족도, 위치, 분위기, 서비스, 목적, 대기시간 (Translation: Food, Food Quantity, Taste, Satisfaction, Location, Atmosphere, Service, Purpose, Waiting Time)
Table 3: Unsupervised Aspect Category Lists for Each Domain
As shown in Table 3, the final number of aspect categories was determined to be 7 for the hotel domain, 9 for the food domain, and 10 for the beauty domain.
5.3 Multi-Label Annotation of Review Data
Fig. 3 illustrates the overall overview of the proposed MUSCAD  framework.
Figure 3: Pipeline of MUSCAD : Aspect Term Extraction and Category Labeling from Reviews
The proposed framework processes input reviews to perform optimal clustering and automatic labeling of aspect terms and aspect categories. Using the constructed dictionaries of aspect categories and aspect terms, we conducted multi-label annotation on new review data. Considering that a single review may correspond to multiple aspect categories, the framework is designed to assign multiple labels to each review. For example, Fig. 3 illustrates that a single review can be assigned to three aspect categories: “Purpose,” “View,” and “Breakfast”. As a result, the automatically labeled dataset consists of 255,110 reviews for the beauty domain, 263,395 for the hotel domain, and 210,272 for the food domain (see Table 2). This multi-labeled dataset clearly classifies the aspects addressed in each review and serves as a valuable resource for future research such as Aspect-Based Sentiment Analysis and recommendation system development. Examples of the data labeled with multiple aspects are presented in  C.
5.4 Multi-label Classification Results
The Micro-F1 and Macro-F1 scores for each domain are presented in Table 4 and Table 5, respectively.
Model Micro-F1 Macro-F1
BERT-base 97.1 96.3
BERT-large 97.1 96.7
RoBERTa 98.2 96.0
DeBERta-Base 97.3 97.2
Multilingual BERT 98.0 97.5
XLM-RoBERTa 98.4 97.4
Table 4: Multi-Label Classification Results on the English Beauty Aspect-Labeled Dataset
Model Hotel Food
Micro-F1 Macro-F1 Micro-F1 Macro-F1
KLUE-BERT 98.2 98.1 97.6 97.2
KoELECTRA 97.5 97.3 96.7 96.5
KR-BERT 98.3 98.3 98.5 98.3
Multilingual BERT 98.0 97.5 96.6 96.4
XLM-RoBERTa 98.4 98.1 97.3 97.1
Table 5: Multi-Label Classification Results on the Korean Hotel & Food Aspect-Labeled Dataset
In the beauty dataset, Multilingual BERT and XLM-RoBERTa demonstrated the best performance. In particular, XLM-RoBERTa achieved the highest Micro-F1 score of 98.4, while Multilingual BERT recorded a Macro-F1 score of 97.5, indicating relatively balanced performance. In the hotel and food datasets, KR-BERT and XLM-RoBERTa showed strong performance. For the hotel dataset, KR-BERT achieved the highest Macro-F1 score of 98.3, whereas XLM-RoBERTa recorded the highest Micro-F1 score of 98.4. In the food dataset, KR-BERT achieved the highest performance, with a Micro-F1 score of 98.5 and a Macro-F1 score of 98.3. Overall, the experimental results confirm that various pre-trained language models perform well on the multi-aspect labeling dataset. In particular, XLM-RoBERTa consistently achieved high performance across all datasets, while KR-BERT demonstrated strong competitiveness in the Korean datasets. These results verify that the labels automatically generated by the framework are of sufficient quality for training supervised models and that various pre-trained language models can achieve excellent performance on the proposed dataset.
5.5 Comparison of Multi-Label Classification and LLM Results
To evaluate the performance of MUSCAD, we conducted comparative experiments with large language models (LLMs). This was done to assess the feasibility of LLM-based labeling and to verify how consistently and accurately fine-tuned models can perform labeling. We extracted 1,000 samples per domain for evaluation. Table 6 and Table 7 compare the results of multi-aspect labeling performed by fine-tuned models and LLMs (GPT-3.5-turbo and GPT-4o-mini) for the beauty, hotel, and food domains, respectively. The LLM-based labeling was conducted using prompts under zero-shot and few-shot settings, and the prompts are provided in D.
Model Micro-F1 Macro-F1
Fine-tuned (after MUSCAD)
BERT-base 74.3 53.0
BERT-large 77.5 64.9
RoBERTa 75.0 58.0
DeBERta-Base 66.4 43.4
Multilingual BERT 70.9 49.9
XLM-RoBERTa 71.4 47.1
LLM (Zero-shot)
ChatGPT (GPT-3.5-turbo) 63.6 54.4
GPT-4o-mini 52.3 49.7
LLM (Few-shot)
ChatGPT (GPT-3.5-turbo) 57.2 58.9
GPT-4o-mini 55.7 49.4
Table 6: Comparison Between Multi-Label Classification and LLM Results on the English Beauty Aspect-Labeled Dataset, evaluated on a test set of 1,000 samples
Model Hotel Food
Micro-F1 Macro-F1 Micro-F1 Macro-F1
Fine-tuned (after MUSCAD)
KLUE-BERT 86.0 85.4 76.1 75.6
KoELECTRA 76.7 67.1 61.3 60.2
KR-BERT 87.2 84.8 74.1 72.2
Multilingual BERT 81.9 78.2 55.8 50.5
XLM-RoBERTa 61.6 48.8 71.1 69.2
LLM (Zero-shot)
ChatGPT (GPT-3.5-turbo) 71.1 70.0 62.3 50.5
GPT-4o-mini 73.2 70.5 57.4 47.7
LLM (Few-shot)
ChatGPT (GPT-3.5-turbo) 73.2 65.4 61.0 50.7
GPT-4o-mini 76.0 72.7 61.1 48.6
Table 7: Comparison Between Multi-Label Classification and LLM Results on the Korean Hotel & Food Aspect-Labeled Dataset, evaluated on text set of 1,000 samples
The comparison results can be summarized in three key findings.
 Performance  Fine-tuned models performed labeling based on data trained for specific domains, and as a result, generally achieved higher Micro-F1 and Macro-F1 scores. In particular, KR-BERT and XLM-RoBERTa showed the best performance in the hotel and food domains, while BERT-large and RoBERTa performed relatively well in the beauty domain. On the other hand, the labeling results using LLMs were generally inferior to those of the fine-tuned models. Although the LLMs demonstrated the ability to interpret sentence meaning and assign appropriate aspects, issues were observed such as assigning incorrect labels or producing inconsistent results for the same sentence (see Table 8).
Review Text MUSCAD LLM-Based Labels
Two men traveled, and perhaps because it was off-season, there were hardly any people; it was quiet and offered good value for the price. Satisfaction Purpose Satisfaction Location
The tangsuyuk is also tasty, but since it’s located on the first floor of a building where many companies are situated, there’s a really long wait at lunchtime. Food Satisfaction Waiting Time Taste Location Waiting Time
I returned this item because it had terrible coverage, similar to using a skin tone primer Color Satisfaction () Satisfaction
Table 8: Comparison of Aspect Category Labeling by MUSCAD and an LLM-Based Model. A green check () indicates a correctly identified aspect, whereas a red cross () denotes an incorrect or missing aspect.
As shown in Table 8, for the same review, the fine-tuned model correctly predicted the aspects “Satisfaction” and “Purpose” whereas the LLM assigned “Satisfaction” and “Location” resulting in partially inaccurate labeling. This suggests that while LLMs are capable of interpreting context and extracting appropriate aspects, they tend to assign semantically inappropriate labels in certain contexts.
Reproducibility  LLM-based labeling exhibited issues with reproducibility. When the same review was input multiple times, the resulting labels were often inconsistent, especially in few-shot settings. This indicates that the outputs of LLMs can vary significantly depending on prompt engineering and the structure of input examples, suggesting that fine-tuned models may be a more reliable alternative for tasks where consistency is critical, such as large-scale data labeling.
Cost  Multi-aspect labeling using LLMs has the advantage of being applicable to various domains without additional data training. However, when processing large volumes of data, the cost becomes a significant drawback. With models such as GPT-4o-mini and GPT-3.5-turbo, API call costs accumulate rapidly, and labeling hundreds of thousands of reviews can lead to substantial expenses. In contrast, fine-tuned models require initial training costs, but once trained, they can efficiently process large datasets without additional costs. Moreover, they can be run on local servers, enabling real-time labeling without delays from API calls. In this regard, fine-tuned models offer better cost-efficiency and scalability compared to LLMs.
In conclusion, while LLM-based multi-aspect labeling can be useful for quickly labeling small datasets, fine-tuned models prove to be a more practical solution for large-scale data processing. They demonstrate higher reproducibility and more consistent performance compared to LLMs, and are more cost-effective, making them better suited for real-world applications.
5.6 Human Evaluation
To qualitatively evaluate the quality of the automated MUSCAD framework, we conducted a Human Evaluation. To assess whether the labels generated by MUSCAD are semantically appropriate in practice, three domain experts were recruited to manually perform multi-aspect category labeling on 100 review samples from each domain. Clear guidelines explaining each aspect category and the labeling criteria were provided to the evaluators to ensure consistency, and the guideline is illustrated in Fig. 4.
Human Evaluation Criteria This survey is a Human Evaluation task that involves selecting appropriate labels based on review data. Please read the given reviews carefully and choose the most relevant label (s). Each item focuses on describing a specific aspect of the review. The reviews are from Amazon’s Beauty & Personal Care product category. The labeling method and guidelines are well explained below, so please read them thoroughly before proceeding. Labeling Method: Select the most relevant label (s) based on the content of the review. If multiple labels apply, check all applicable labels. Labeling Guidelines: Satisfaction: Expressions related to overall satisfaction with the experience. Ingredients: Mentions of product ingredients, raw materials, or composition. Usage Method: References to product usage methods or applications (e.g., face, body, cosmetics, etc.). Improvement: Aspects that have improved after using the product. Color: Opinions related to the product’s color. Hair: Effects on hair after using the product or changes related to hair condition. Packaging: Mentions of packaging design, durability, shipping condition, etc. Scent: Descriptions of fragrance, intensity, longevity, or types of scent. Purchase: Mentions of purchasing experience, reasons for purchase, or target audience for the product.
Figure 4: Human Evaluation Criteria for Qualitative Assessment
The evaluation was conducted by comparing the labels generated by MUSCAD with those assigned by each evaluator using Micro-F1 and Macro-F1 scores. The results are presented in Table 9.
Evaluator Hotel Food Beauty
Micro-F1 Macro-F1 Micro-F1 Macro-F1 Micro-F1 Macro-F1
Evaluator 1 78.6 79.3 60.1 61.1 67.7 61.2
Evaluator 2 69.4 72.5 74.1 71.7 66.7 61.6
Evaluator 3 69.1 72.1 75.8 72.9 69.7 61.6
Mean (Std) 72.4 (4.4) 74.6 (3.4) 70.0 (7.0) 68.6 (5.4) 68.0 (1.2) 61.5 (0.2)
Table 9: Human Evaluation Results with Average, Standard Deviation, and MUSCAD Comparison
According to the Human Evaluation results presented in Table 9, the labels automatically generated by MUSCAD maintained a consistent level of performance when compared to those assigned by human evaluators. This indicates that the proposed framework performs aspect extraction and classification tasks at a reliable level within specific domains. The Human Evaluation also revealed differences among evaluators, with some domains showing inconsistent application of labeling criteria. This suggests that human interpretation may vary depending on the individual, and that a clear understanding of the guidelines is essential for maintaining consistency in labeling specific aspect categories. In particular, there were cases where the same review was labeled differently by different evaluators, highlighting the potential for subjectivity in manual labeling. In contrast, MUSCAD applies automated labeling based on a consistent standard, which minimizes variability among evaluators and produces more uniform results. Manual labeling requires expert involvement and is both time-consuming and costly. Even when guidelines are provided, differences in interpretation can arise. However, automated labeling using the proposed framework overcomes these limitations by enabling fast and consistent processing of large-scale data.
5.7 Summary
As demonstrated through various experiments in Section 5, the multi-aspect labels generated by the proposed framework are of sufficient quality for use in model training. The multi-label classification experiments using various fine-tuned models yielded high Micro-F1 and Macro-F1 scores, indicating that the automatically labeled data is suitable for supervised learning. Furthermore, we compared the performance of fine-tuned models with few-shot labeling using GPT-3.5-turbo and GPT-4o-mini to evaluate the capabilities of LLMs in multi-aspect labeling. The results showed that while LLMs were able to assign appropriate labels in some cases, their overall performance was inferior to that of the fine-tuned models. LLMs also exhibited issues with reproducibility and incurred higher costs. In addition, Human Evaluation confirmed that the automatic labels produced by MUSCAD achieved a level of reliability and consistency comparable to manual labeling by experts, and were more effective in reducing subjectivity among evaluators. These findings suggest that the proposed framework offers a more practical solution for applying multi-aspect labeling to large-scale review data.
6 Discussion
6.1 Conclusion
In this study, we proposed MUSCAD, an unsupervised multi-aspect labeling framework designed to automatically assign aspect categories to multilingual and multi-domain review data. While prior research has largely relied on supervised learning approaches that require large amounts of manually labeled data, our framework integrates K-means–based aspect candidate generation, contextual representation learning using Multi-Head Attention, and classification optimization via Max-Margin Loss, enabling multi-aspect labeling without predefined vocabularies. Experimental results demonstrated that fine-tuned classification models trained on the resulting multi-aspect labeling dataset achieved high F1-scores, validating the reliability of the automatically generated labels. Compared to large language models (LLMs), MUSCAD provided more consistent and reproducible labeling outcomes. In addition, Human Evaluation confirmed that the quality of the automatically generated labels was comparable to that of manual labeling by experts. Finally, the MUSCAD  framework has shown scalability and practical applicability in real-world environments through large-scale automatic labeling across multiple languages and domains.
6.2 Future Work
This study proposed and evaluated the performance of the unsupervised MUSCAD framework across diverse domains. However, several directions remain for future improvements and extensions. First, we aim to build a fully automated labeling system using AI agents. Currently, partial manual intervention is required for matching aspect categories. In future work, we plan to eliminate manual processes entirely by incorporating AI agents and establishing an automated feedback loop, ultimately enabling fully autonomous labeling. Second, developing aspect-aware review summarization techniques is another key direction. While this study focuses on identifying the aspects discussed in each review, it does not provide functionality for summarizing the core content of individual reviews. As a next step, we aim to build a model that generates aspect-level summaries using the labeled data. For example, in hotel reviews, aspects such as “Room”, “Location” and “Service” can be automatically extracted, and the content related to each aspect can be summarized and presented to users. This would help enhance user experience and improve the utility of review data. Building upon these directions, we intend to further refine MUSCAD and develop an automated labeling and summarization system that can be applied in real-world settings. By integrating a fully automated labeling pipeline based on AI agents with aspect-aware review summarization, we envision the creation of a more efficient and scalable NLP-based review analysis system.
Appendix A Experiment on the Number of Words per Aspect Category
The number of words extracted for each aspect category affects both the model’s performance and the data coverage. To determine the optimal number of words, we conducted experiments under various settings to compare performance. Specifically, we constructed multi-aspect labeling datasets by extracting the top 50, 100, 150, 200, and 250 words for each aspect category, and evaluated the performance using the XLM-RoBERTa model. The main objective of this experiment was to analyze how the number of extracted words influences classification performance (F1-score) and coverage (i.e., the proportion of relevant words included within each aspect category). Through this experiment, we aimed to identify the optimal number of words that maximizes generalization performance while maintaining model accuracy. The results and analysis are presented in Table A.1 and Table A.2.
Word Count Hotel Food Beauty
50 97.17% 76.14% 73.32%
100 98.41% (+1.24%) 87.55% (+11.41%) 85.43% (+12.11%)
150 98.60% (+0.19%) 84.89% (-2.66%) 92.24% (+6.81%)
200 98.81% (+0.21%) 83.85% (-1.04%) 95.20% (+2.96%)
250 98.96% (+0.15%) 97.51% (+13.66%) 96.31% (+1.11%)
Table A.1: Coverage of Top-N Words across Three Domains (Hotel, Food, Beauty). Each row corresponds to a specific number of words (e.g., 50, 100, 150, etc.), and the table shows the cumulative percentage of tokens covered in each domain. Values in parentheses indicate the relative change in coverage compared to the immediately preceding row.
As shown in Table A.1, increasing the number of words per aspect category from 50 to 100 led to a significant improvement in coverage across all domains. In particular, the food and beauty domains showed substantial increases of +11.41% and +12.11%, respectively. At 150 words, coverage continued to improve in the hotel and beauty domains; however, a slight decrease of –2.66% was observed in the food domain. When increasing the word count beyond 200, the gains became marginal or even negative, suggesting that including excessive terms may lead to performance degradation. Based on these findings, and considering both coverage improvement and generalization performance, we selected 150 words as the optimal number for each aspect category.
A.1 Performance Analysis by Number of Words
  Hotel Food Beauty Word Count Micro-F1 Macro-F1 Micro-F1 Macro-F1 Micro-F1 Macro-F1 50 0.9785 0.974 0.9405 0.9022 0.9655 0.8658 100 0.9731 (+0.54%) 0.9726 (+0.14%) 0.9291 (-1.14%) 0.9102 (+0.8%) 0.9477 (-1.78%) 0.8774 (+1.16%) 150 0.9751 (+0.20%) 0.9732 (+0.06%) 0.9097 (-1.95%) 0.8965 (-1.37%) 0.9439 (-0.38%) 0.8841 (+0.67%) 200 0.9778 (+0.27%) 0.9757 (+0.25%) 0.9051 (-0.46%) 0.8853 (-1.12%) 0.9429 (-0.10%) 0.8686 (+1.55%) 250 0.9713 (-0.65%) 0.9688 (-0.69%) 0.9202 (+1.51%) 0.9066 (+2.13%) 0.9412 (-0.17%) 0.8708 (+0.22%)
Table A.2: Performance and Coverage of XLM-RoBERTa with Varying Top-N Word Counts. The table reports the Micro-F1 and Macro-F1 scores for Hotel, Food, and Beauty domains when selecting the top 50, 100, 150, 200, and 250 words, with the values in parentheses indicating the relative change compared to the previous setting.
Table A.2 presents the classification performance of XLM-RoBERTa according to the number of extracted words. When using 50 words, both Micro-F1 and Macro-F1 scores were relatively low across all domains. When the number of words exceeded 200, a performance drop was observed in some domains. Using 150 words proved to be the optimal setting, as it maintained strong performance in the hotel and beauty domains while minimizing the performance degradation in the food domain. Qualitative analysis also confirmed that the 150-word configuration produced the most balanced labeling results and helped prevent performance degradation caused by the inclusion of unnecessary terms. Based on these findings, we selected 150 words as the final optimal number for each aspect category.
Aspect Category Candidates (Food Domain) Taste, Satisfaction, Service, Food Quantity, Menu, Waiting Time, Atmosphere, Cleanliness, Price, Purpose, Delivery, Location, Amenities
Aspect Category Candidates (Beauty Domain) Pigmentation, Persistence, Moisture, Skin Type, Spreadability, Usage Method, Scent, Color, Seasonal Use, Irritation, Ingredients, Hair, Improvement, Satisfaction, Packaging, Purchase
Figure A.1: Food & Beauty Aspect Category Candidates. The figure presents candidate aspect categories for the Food and Beauty domains, which are used for prompt-based assignment in our framework.
Appendix B Aspect Category & Aspect Term
This appendix presents the aspect categories selected in this study along with the corresponding lists of aspect terms. The aspect categories and associated aspect terms for the hotel, food, and beauty domains are provided in Table B.1, Table B.2, and Table B.3, respectively.
Aspect Category Aspect Term
룸 (Room) 세면대, 드라이기, 슬리퍼, 침실, 소파, 화장실, 어메니티, 커피포트, 칫솔, 조명, 면봉, 로션 (Washbasin, Hairdryer, Slippers, Bedroom, Sofa, Bathroom, Amenity, Coffee Pot, Toothbrush, Lighting, Cotton Swab, Lotion)
목적 (Purpose) 여자친구, 남자친구, 엄마, 신혼여행, 휴가, 부모님, 우정, 여행, 호캉스, 연인, 커플, 기념, 신혼, 결혼기념일 (Girlfriend, Boyfriend, Mom, Honeymoon, Vacation, Parents, Friendship, Travel, Hotel Stay, Lover, Couple, Anniversary, Newlywed, Wedding Anniversary)
위치 (Location) 공항버스, 재래시장, 이마트, 협재, 횟집, 해운대, 국제공항, 수산시장, 해수욕장, 미술관, 터미널, 관광지(Airport Bus, Traditional Market, Emart, Hyeopjae, Seafood Restaurant, Haeundae, International Airport, Fish Market, Beach, Art Museum, Terminal, Tourist Spot)
부대시설 (Facilities) 레스토랑, 생맥주, 베이커리, 다이닝, 디너, 자판기, 뷔페, 주차장, 치킨, 무한리필, 한식, 스프, 미니바 (Restaurant, Draft Beer, Bakery, Dining, Dinner, Vending Machine, Buffet, Parking Lot, Chicken, Unlimited Refill, Korean Cuisine, Soup, Mini-bar)
뷰 (View) 일몰, 풍경, 전경, 야경, 노을, 항구, 정원, 한라산, 해돋이, 감상, 바다, 전망, 뷰, 햇살, 오션(Sunset, Scenery, Panorama, Night View, Afterglow, Harbor, Garden, Hallasan, Sunrise, Appreciation, Sea, View, Sunshine, Ocean)
서비스 (Service) 물어보다, 죄송하다, 말씀, 대처, 상담, 말씀드리다, 전화하다, 카운터, 요청, 프론트, 응대, 감사 (Inquire, Apologize, Communicate, Handle, Consult, Inform, Call, Counter, Request, Front Desk, Respond, Appreciate)
만족도 (Satisfaction) 평범하다, 오래되다, 일반, 그럭저럭, 어수선하다, 노후하다, 협소하다, 양호하다, 퀄리티, 나쁘다, 열악하다, 뛰어나다, 럭셔리, 훌륭하다 (Average, Old, General, So-so, Disorganized, Outdated, Cramped, Good, Quality, Bad, Poor, Excellent, Luxurious, Superb)
Table B.1: Hotel Aspect Categories and Representative Terms (Korean with English Translation). These categories and representative terms are automatically derived by our framework from hotel review texts, and the Korean aspect terms are presented along with their corresponding English translations.
Aspect category Aspect term
음식 (Food) 포테이토, 베이컨, 머쉬룸, 발사믹, 고르곤졸라, 프로슈토, 에그베네딕트, 칼국수, 보쌈, 된장찌개 (Potato, Bacon, Mushroom, Balsamic, Gorgonzola, Prosciutto, Egg Benedict, Knife-cut Noodles, Bossam, Soybean Paste Stew)
음식량 (Food Quantity) 골고루, 여러가지, 세트, 글라스, 사이드, 무제한, 혼합, 플래터, 단일, 스몰, 메인, 글라스, 디시, 옵션 (Evenly, Various, Set, Glass, Side, Unlimited, Mixed, Platter, Single, Small, Main, Glass, Dish, Option)
맛 (Taste) 고소하다, 단맛, 조화롭다, 짭짤하다, 산뜻하다, 부드러움, 느끼하다, 달콤하다, 씁쓸하다 (Savory, Sweetness, Harmonious, Salty, Refreshing, Softness, Greasy, Sweet, Bitter)
만족도 (Satisfaction) 부실하다, 비싸다, 가격대비, 평범하다, 부족하다, 애매하다, 아쉬움, 실망하다, 저렴하다, 심각하다 (Inadequate, Expensive, Price-to-Value, Ordinary, Lacking, Vague, Regret, Disappointed, Affordable, Serious)
위치 (Location) 뚝섬역, 이대, 안국역, 홍대, 신촌, 성수동, 서울, 강남, 뚝도, 대학가, 을지로, 이태원 (Ttukseom Station, Ewha (Edae), Anguk Station, Hongdae, Sinchon, Seongsu-dong, Seoul, Gangnam, Ttukdo, University Area, Euljiro, Itaewon)
분위기 (Atmosphere) 아기자기하다, 화려하다, 예쁘다, 화사하다, 세련되다, 고급, 모던, 빈티지, 심플, 힙하다, 엔틱(Charming, Glamorous, Pretty, Bright, Sophisticated, Luxurious, Modern, Vintage, Simple, Hip, Antique)
서비스 (Service) 말투, 표정, 아주머니, 서빙, 태도, 죄송하다, 종업원, 알바, 매니저, 응대, 불친절, 불쾌하다 (Tone of speech, Facial expression, Middle-aged woman, Serving, Attitude, Apologize, Staff, Part-timer, Manager, Customer service, Unfriendly, Unpleasant)
목적 (Purpose) 남자친구, 남편, 와이프, 엄마, 아빠, 부모님, 동료, 결혼기념일, 신년, 가족, 생일, 외식, 명절 (Boyfriend, Husband, Wife, Mom, Dad, Parents, Colleague, Wedding Anniversary, New Year, Family, Birthday, Dining Out, Holiday)
대기시간 (Waiting Time) 50분, 30분, 10분, 웨이팅, 기다림, 브레이크, 테이블링, 라스트, 오픈런, 평일 (50 minutes, 30 minutes, 10 minutes, Waiting, Waiting time, Break, Tabling, Last order, Open run, Weekday)
Table B.2: Food Aspect Categories and Representative Terms (Korean with English Translation). These categories and representative terms are automatically derived by our framework from food review texts, and the Korean aspect terms are presented alongside their corresponding English translations.
Aspect category Aspect term
Improvement dehydrated, dryness, calm, aging, improved, plump, tigh, acneprone, smoothing, redness, youthful, hydrate, nourish, wrinkle, oiliness, glowing
Ingredients tocopheryl, ester, acetate, tocopherol, sativa, helianthus, palmitate, glycine, kernel, hydrogenated, panthenol, silica, annuus, argania, allantoin
Satisfaction actually, difficult, okay, hard, harder, good, well, great, nice, kind, tend, better, stronger, normal, properly, terrible, regular, exactly, strange, awesome, entirely
Color shimmer, highlighter, nude, beige, flittery, neon, coral, darker, bright, pink, gold, brown, bronze, rosy, color, yellow, blue, blond, gray, silver, smokey, navy, ashy
Hair straightener, wavy, tail, wave, curly, bun, bang, pony, curl, tangle, anchor, wig, shear, clip, tapered, turban, mane, dryer, heated, hair, shaft, trim, pin, snag
Packaging storage, sippered, tray, pouch, sleeve, soered, insert, zip, bag, tin, pocket, sealed, square, closure, foil, removable, encased, freebie, bowl, hook, screw, hole, ribbon
Scent pungent, fruity, flowery, musky, woodsy, spicy, floral, earthy, minty, sweet, citrusy, patchouli, sandalwood, vanilla, lavender, peppermint, tabacco, bergamot, jasmine, cedarwood
Usage method cleanse, pat, toner, massage, wipe, patted, rub, scrub, soak, showering, dab, rinsing, apply, wiped, exfoliate, spread, scrubbing, lather, absorbed, distribute, rinsed, lathering
Purchase bought, ordered, shopping, paid, acquired, picked, gifted, returned, exchanged, mom, sister, teen, son, dad, female, family, boyfriend, husband, kid, lady, boy, oldest, mail
Persistence lasted, extended, duration, progress, multiple, continuous, third, second, minimum, maximum, later, final, broke, fourth, third, repeated
Table B.3: Beauty Aspect Categories and Representative Terms (Korean with English Translation). These categories and representative terms are automatically derived by our framework from beauty review texts, with the Korean aspect terms presented alongside their corresponding English translations.
Appendix C Multi-Aspect Category Labeling Results
This appendix presents the prompts used for multi-aspect labeling with LLMs in the comparative experiments described in Section 5.5. In this study, we designed appropriate prompts to evaluate whether LLMs can assign correct aspect categories based on review sentences under zero-shot and few-shot settings. The prompts are shown in Fig. D.1, Fig. D.2, Fig. D.3, Fig. D.4, and Fig. D.5.
Review Multi-Aspect Category
제주시내에 위치하고 있어서 편리하게 이용하기 좋아요 (It is conveniently located in Jeju City, making it easy to use.) 만족도, 위치 (Satisfaction, Location)
아침일출과 밤바다 보이는 바닷가 객실을 추천합니다 (I recommend the beachside room with views of the sunrise and the night sea.) (View)
숙소 수영장 조식까지 최고였어요 (The accommodation, swimming pool, and breakfast were excellent.) 만족도, 부대시설 (Satisfaction, Facilities)
아이들을 위한 키즈카페와 다양한 부대시설 너무 맘에든 첫 숙소였어요 (This was our first hotel that we really liked because of its kids’ cafe and various facilities for children.) 목적, 부대시설 (Purpose, Facilities)
여자친구와 500일 기념 및 생일 기념으로 호텔에 방문했습니다 (I visited the hotel to celebrate 500 days with my girlfriend and also for her birthday.) 목적 (Purpose)
넓기도 하고 욕조 또한 정말 좋습니다 (It is spacious and the bathtub is also excellent.) 만족도, 룸 (Satisfaction, Room)
위치는 올래시장과 이중섭거리 바로 옆이라 뭐 먹기는 최고임 (The location is excellent for dining as it is right next to Ollae Market and Lee Jung-seop Street.) 만족도, 위치 (Satisfaction, Location)
깨끗하고 친절한 직원들에 감동입니다 (I am impressed by the cleanliness and the friendly staff.) 룸, 서비스 (Room, Service)
피트니스센터 왜케 기구들 너무 좋나요 (The gym equipment at the fitness center is amazing.) 만족도, 부대시설 (Satisfaction, Facilities)
침대 더블 1개 싱글 2개 배역 어른 둘 초딩 하나 가족인데 각자 하나씩 편하게 잤어요 (We had one double bed and two single beds – perfect for our family of two adults and one child, as everyone slept comfortably.) 만족도, 룸, 목적 (Satisfaction, Room, Purpose)
Table C.1: Examples of Hotel Reviews with Multi-Aspect Category Labels (Korean and English). Each review is presented in Korean alongside its English translation, and the corresponding multi-aspect category labels are automatically assigned by our framework. This table illustrates how a single review can be associated with multiple categories (e.g., Satisfaction, Location, Facilities), reflecting various aspects of the hotel experience.
Review Multi-Aspect Category
제 입맛에는 딤섬 특히 맛났고 광동식 탕수육 좋았어요 (The dim sum was especially delicious to my taste, and the Cantonese-style sweet and sour pork was great.) 음식, 맛 (Food, Taste)
프라이빗한 모임으로 제격이네요 (Perfect for a private gathering.) 목적 (Purpose)
창덕궁역 뷰맛집 엄마랑 호캉스 여행기념으로 다녀왔어요 (A great restaurant with a view near Changdeokgung Station. I visited to celebrate a hotel staycation with my mom.) 위치, 분위기, 목적 (Location, Atmosphere, Purpose)
연말 단체 모임 갔었는데 좋았어요. 양은 적음 (I went there for a year-end group gathering, and it was nice. However, the portions were small.) 만족도, 목적, 음식량 (Satisfaction, Purpose, Food Quantity)
애견동반되서 강아지와 행복한 시간 보냈네요 (I was able to bring my dog and had a great time together.) 목적 (Purpose)
밀크티 정말 맛있고 패션후르츠 음료도 아주 맛있습니다 (The milk tea was really delicious, and the passion fruit drink was also very tasty.) 음식 (Food)
그런데도 예의 없이 손님에게 화내고 소리지르는 모습에 화가 나 리뷰를 작성합니다 (However, I was angered to see the staff rudely yelling at customers, which prompted me to write this review.) 서비스 (Service)
웨이팅이 조금 있었지만 금방 빠졌어요 (There was a bit of a wait, but it cleared up quickly.) 대기시간 (Waiting Time)
부모님도 좋아하시고 메뉴가 어른들 드시기에도 좋습니다 (My parents liked it, and the menu is well-suited for older guests.) 만족도, 목적 (Satisfaction, Purpose)
커피랑 젤라또 전문이라고 해서 들어왔는데 전문답게 맛있어요 (I came in because they specialize in coffee and gelato, and it truly tasted professional.) 음식 (Food)
Table C.2: Examples of Food Reviews with Multi-Aspect Category Labels (Korean and English). Each review is displayed in Korean along with its English translation, and the corresponding multi-aspect category labels are automatically assigned by our framework. This table illustrates how a single review can span multiple categories (e.g., Food, Taste, Location, Purpose), reflecting various facets of the dining experience.
Review Multi-aspect category
These are extremely dull and will wreck your nails. Improvement
And the dark brown is not brown, it’s rust/copper. Color
Works wonders. They have other scents as well and I like all of them, but the classic is my favorite. Scent, Satisfaction
I found them on Amazon and bought a 3 pack. They are compact and easy to carry. Satisfaction, Purchase, Packaging
Wonderful bath salt - smells great. Scent, Satisfaction, Usage Method
After I washed and dried my hair, I lifted parts and gave them a quick squirt, let them dry, combed them down a little and did a light spray over the rest of my hair. Hair, Usage Method
This complex is great and contains all you need to combat aging and fine lines. Ingredients, Satisfaction, Improvement
I did apply a small dab of shea butter to that spot and it no longer itches AT ALL. Satisfaction, Usage Method, Improvement
The rose quartz is a very light pink and white… very pretty. Usage Method, Color
It’s been about 5 weeks and my cracks are almost fully healed and even my pedicurist was shocked! Persistence, Improvement
Table C.3: Examples of Beauty Reviews with Multi-Aspect Category Labels (English). Each review is presented in English, and the corresponding multi-aspect category labels are automatically assigned by our framework. This table demonstrates how a single review can reflect various aspects (e.g., color, scent, satisfaction, improvement), offering insight into diverse beauty product experiences.
Appendix D Prompt
This appendix presents the prompts used for multi-aspect labeling with LLMs in the comparative experiment described in Section 5.5. In this study, we designed appropriate prompts to evaluate whether LLMs can assign correct aspect categories to review sentences under zero-shot and few-shot settings. The prompts are shown in Fig. D.1, Fig. D.2, Fig. D.3, Fig. D.4, and Fig. D.5.
Zero-shot Prompt (Hotel Domain) 당신은 주어진 리뷰 문장을 읽고 아래의 aspect 중 하나 이상으로 분류하는 작업을 맡았습니다.
반드시 아래 aspect 목록에서만 선택하고, 제공되지 않은 키워드는 절대 포함하지 마세요.
Aspect 목록: 목적, 룸, 위치, 부대시설, 만족도, 뷰, 서비스 Instruction: 리뷰를 가장 잘 설명하는 하나 이상의 aspect를 선택하세요. 문맥과 세부 사항을 주의 깊게 읽고 적합한 측면을 판단하세요. 선택한 aspect 키워드만 쉼표로 구분하여 반환하세요. 설명이나 추가 텍스트는 포함하지 마세요. Labeling guideline: 목적: 숙소 방문 목적 (예: 신혼여행, 가족여행 등)과 관련된 표현 룸: 객실의 청결, 방음, 어메니티 등 객실 환경과 관련된 표현 위치: 숙소의 위치, 접근성, 주변 시설과의 거리 관련 표현 부대시설: 수영장, 조식 등 숙소 내 제공되는 부대시설과 관련된 표현 만족도: 숙박 전반에 대한 만족도 또는 가격 대비 가치 관련 표현 뷰: 창문, 발코니 등에서 보이는 전망과 관련된 표현 서비스: 직원의 친절도, 응대 태도, 서비스 품질 관련 표현
You are tasked with categorizing the given review sentence into one or more of the aspects listed below.
You must select only from the provided aspect list and must not include any keywords that are not listed.
Aspect List: Purpose, Room, Location, Facilities, Satisfaction, View, Service Instruction: Select one or more aspects that best describe the review. Carefully read the context and details to determine the most appropriate aspect(s). Return only the selected aspect keywords, separated by commas. Do not include explanations or additional text. Labeling guideline: Purpose: Expressions related to the reason for visiting the accommodation (e.g., honeymoon, family trip) Room: Expressions related to room conditions such as cleanliness, soundproofing, and amenities Location: Expressions related to the accommodation’s location, accessibility, and proximity to nearby facilities Facilities: Expressions related to in-house amenities such as swimming pools and breakfast Satisfaction: Expressions indicating overall satisfaction with the stay or value for money View: Expressions related to the scenery from windows, balconies, etc. Service: Expressions related to staff friendliness, customer service, and service quality
Figure D.1: Zero-shot Prompt for Hotel Aspect Category Labeling (Korean with English Translation).This figure shows a zero-shot prompt used by the LLM to categorize hotel-related review sentences into predefined aspects (e.g., purpose, room, location, facilities, satisfaction, view, service). The original instructions and labeling guidelines are in Korean, with an English translation provided for clarity.
Zero-shot Prompt (Food Domain) 당신은 주어진 리뷰 문장을 읽고 아래의 aspect 중 하나 이상으로 분류하는 작업을 맡았습니다.
반드시 아래 aspect 목록에서만 선택하고, 제공되지 않은 키워드는 절대 포함하지 마세요.
Aspect 목록: 만족도, 맛, 목적, 분위기, 서비스, 위치, 음식, 음식량, 대기시간 Instruction: 리뷰를 가장 잘 설명하는 하나 이상의 aspect를 선택하세요. 문맥과 세부 사항을 주의 깊게 읽고 적합한 측면을 판단하세요. 선택한 aspect 키워드만 쉼표로 구분하여 반환하세요. 설명이나 추가 텍스트는 포함하지 마세요. Labeling guideline: 목적: 방문한 이유 또는 목적 관련 표현 분위기: 인테리어, 조명, 뷰 등 분위기 관련 표현 위치: 장소의 위치, 접근성, 교통 관련 표현 만족도: 전반적인 경험에 대한 만족도 표현 포함, 가성비 포함 음식: 음식의 종류, 품질, 음식명 관련 표현 포함 음식량: 음식의 양, 주문 양과 관련된 표현 대기시간: 대기시간, 줄 서는 시간 관련 표현 맛: 음식의 구체적인 맛과 관련된 표현 서비스: 직원의 태도, 응대 속도, 친절도 및 전반적인 서비스 표현
You are tasked with reading a given review sentence and categorizing it into one or more of the aspects below.
Make sure to select only from the provided aspect list and do not include any additional keywords.
Aspect List: Satisfaction, Taste, Purpose, Atmosphere, Service, Location, Food, Portion Size, Waiting Time Instruction: Select one or more aspects that best describe the review. Carefully read the context and details to determine the most appropriate aspect(s). Return only the selected aspect keywords, separated by commas. Do not include explanations or additional text. Labeling guideline: Purpose: Expressions related to the reason for visiting (e.g., special occasion, family gathering) Atmosphere: Expressions related to interior design, lighting, or view Location: Expressions about accessibility, transportation, and convenience of the place Satisfaction: Expressions indicating overall experience satisfaction, including value for money Food: Mentions of food type, quality, or specific dishes Food Quantity: Expressions related to food quantity or portion sizes Waiting Time: Mentions of wait times, queue lengths, or delays in service Taste: Expressions describing the specific taste of food Service: Expressions regarding staff attitude, responsiveness, friendliness, and overall service quality
Figure D.2: Zero-shot Food Prompt (Korean with English Translation). This figure presents a zero-shot prompt used by the LLM to categorize food-related review sentences into predefined aspects (e.g., satisfaction, taste, purpose, atmosphere, service, location, food, portion size, waiting time). The instructions and labeling guidelines are originally written in Korean, with an English translation included for clarity.
Zero-shot Prompt (Beauty Domain) You are tasked with categorizing the given review sentence into one or more of the aspects listed below.
You must select only from the provided aspect list and must not include any keywords that are not listed.
Aspect: Color, Hair, Improvement, Ingredients, Packaging, Persistence, Purchase, Satisfaction, Scent, Usage Method Instructions: Choose one or more aspects from the provided list that best describe the review. Do not include any aspects that are not in the provided list. Only return the aspect keywords separated by commas. Do not include any explanations or additional text. Labeling guideline: Satisfaction: Expressions related to overall satisfaction with the experience. Ingredients: Mentions of product ingredients, raw materials, or composition. Usage Method: References to product usage methods or applications (e.g., face, body, cosmetics, etc.). Improvement: Aspects that have improved after using the product. Color: Opinions related to the product’s color. Hair: Effects on hair after using the product or changes related to hair condition. Packaging: Mentions of packaging design, durability, shipping condition, etc. Scent: Descriptions of fragrance, intensity, longevity, or types of scent. Purchase: Mentions of purchasing experience, reasons for purchase, or target audience for the product.
Figure D.3: Zero-shot Beauty Prompt (English Only). This figure shows a zero-shot prompt used by the LLM to classify beauty-related review sentences into predefined aspects (e.g., color, hair, improvement, packaging, etc.). All instructions, aspect lists, and labeling guidelines are provided in English, and users must select only from the given aspect categories without adding any unlisted keywords.
Few-shot Example Prompt (Hotel Domain) 리뷰: 깨끗하고 방도넓고 함덕해변도 가까워요  →  위치, 룸
The room was clean and spacious, plus Hamdeok Beach was just a short walk away!  →  Location, Room
리뷰: 공항이랑 정말 가깝고 주변에 편의점이나 식당이 많아서 좋았습니다  →  위치, 만족도
Super close to the airport, and there were plenty of convenience stores and restaurants nearby, which was really convenient.  →  Location, Satisfaction
리뷰: 주변에 이마트 스타벅스 맥도날드 등 편의시설 및 식당 많았구요  →  위치
Lots of places around like E-Mart, Starbucks, and McDonald’s really easy to find anything you need!  →  Location
리뷰: 프론트 직원분들도 정말 친절하시고 객실도 청결하고 좋았어요  →  룸, 서비스, 만족도
The front desk staff were super friendly, and the room was spotless and comfy!  →  Room, Service, Satisfaction
리뷰: 화장실에서 락스냄새가 좀 심하긴했지만 방은 크고 넓고 좋아요  →  룸, 만족도
There was a strong bleach smell in the bathroom, but the room itself was big and comfortable.  →  Room, Satisfaction
리뷰: 조식 또한 간단하면서 먹을만해서 만족스럽습니다  →  부대시설, 만족도
Breakfast was simple but decent, and I was happy with it.  →  Facilities, Satisfaction
리뷰: 가족 커플 둘다 추천드려요  →  목적
Great for both families and couples—highly recommend!  →  Purpose
리뷰: 주차장 들어가는 길 하드코어 한거 빼곤 괜찮았어요  →  부대시설
The road leading to the parking lot was a bit rough, but everything else was totally fine.  →  Facilities
리뷰: 카운터 직원분이 너무 친절하고 시설도 다 좋았어요  →  만족도, 서비스
The staff at the counter were super nice, and the facilities were all in great condition.  →  Satisfaction, Service
Few-shot Example Prompt (Food Domain) 리뷰: 카페 인테리어도 이쁘고 2층은 뷰가 이뻤어요  →  분위기
The cafe interior was beautiful, and the view from the second floor was amazing.  →  Atmosphere
리뷰: 냄새도 안나고 고기도 부드럽고 이래저래 좋았습니다  →  맛, 만족도
No bad smell, the meat was tender, and everything was great overall.  →  Taste, Satisfaction
리뷰: 점심시간 줄은 길고 웍이 하나라 요리 나오는 시간이 좀 걸려요  →  만족도, 대기시간
The lunch line is long, and since there’s only one wok, food takes some time to come out.  →  Satisfaction, Waiting Time
리뷰: 서울숲 올 일 있을때마다 들리고 싶네요  →  위치
I’d love to stop by whenever I visit Seoul Forest.  →  Location
리뷰: 버섯 전골 맛있어요  →  음식
The mushroom hot pot is delicious.  →  Food
리뷰: 생일메세지까지 섬세하게 챙겨주셔서 너무 감동받았어요  →  목적, 서비스
They even prepared a birthday message for me, which was so thoughtful and touching.  →  Purpose, Service
리뷰: 내부가 예쁘고 남자직원분 친절해요  →  분위기, 서비스
The interior is pretty, and the male staff member was very friendly.  →  Atmosphere, Service
리뷰: 크림 너무 부드럽고 담백 시트도 촉촉해요  →
The cream is super smooth and light, and the cake layers are moist.  →  Taste
리뷰: 저랑 친구는 그래도 음식 3개를 시켜서 배부르게 먹고 왔습니다  →  만족도, 음식량
My friend and I ordered three dishes and left feeling full and satisfied.  →  Satisfaction, Food Quantity
Figure D.4: Few-Shot Examples for the Hotel & Food Domains. This figure demonstrates how examples of multi-aspect category labels assigned to reviews in the Hotel and Food domains are provided to the model in a few-shot setting. By referring to the aspect categories assigned to each review, the model is guided to classify new sentences more accurately.
Few-shot Example Prompt (Beauty Domain) Review: Opened the package & instant migraine.  →  Improvement Review: Beautiful palette very pleased!  →  Color Review: The price and deal can’t be beat, lasts awhile.  →  Satisfaction Review: My hair is so soft, I didn’t need conditioner.  →  Satisfaction, Improvement Review: Removes makeup and dirt like magic.  →  Satisfaction, Usage Method Review: These socks are utterly amazing!!!  →  Ingredients, Usage Method Review: Worked within a few hours.  →  Persistence Review: Argan oil knows what they are doing!  →  Satisfaction, Ingredients
Figure D.5: Few-Shot Examples for the Beauty Domains. This figure demonstrates how examples of multi-aspect category labels assigned to reviews in the Beauty domains are provided to the model in a few-shot setting. By referring to the aspect categories assigned to each review, the model is guided to classify new sentences more accurately.